The magnetic field strenght of a solenoid is 0.0270T. Its radius is 0.40 m and length is 0.40 m. How many turns are there in the solenoid if the steady current passing through it is 12.0 A?

1
Expert's answer
2022-03-16T18:28:55-0400

The magnetic field inside the solenoid is given as follows:


"B = \mu_0 \dfrac{NI}{L}"

where "I = 12A" is the current, "L = 0.4m" is the length of solenoid, "\mu_0 = 1.26\times 10^{-6}N/A^2" is the magnetic constant, "N" - number of turns. Substituting "B = 0.027T", get:


"N = \dfrac{BL}{\mu_0I} = \dfrac{0.027\cdot 0.4}{1.26\times 10^{-6}\cdot 12} \approx 710turns"

Answer. 710 turns.